OpenStudy (anonymous):

x/10 + y/4 + z/20 = 2.5 x/7.5 + y/6 + z/15 = 3 x/15 + y/3 + z/40 = 1.75 i forget how to clear the fractions?

myininaya (myininaya):

It's look like the first equation. What is the lCM of the denominators? 20? Multiply the first equation by 20. 2nd equation: Note 7.5=15/2. So the denominator of the first term is 15. So Multiplying by 30 will clear your fractions up. 3rd equation: 1.75=1and 3/4=7/4. So multiplying by 120 should do it